OpenStudy (anonymous):

I will medal and fan Can you write twenty-four twenty-firsts as a mixed number? Explain.

OpenStudy (kropot72):

Twenty-one twenty-firsts make one (21/21 = 1). That leaves 3 twenty-firsts over. 3/21 = 1/7. That is why \[\large \frac{24}{21}=1\frac{1}{7}\]

OpenStudy (anonymous):

Ok thank you very much

OpenStudy (kropot72):

You're welcome :)